# Harnessing Large Language Models for Disaster Management: A Survey

### v1

@@ 5 Challenges and Opportunities
Large language models (LLMs) show promise in disaster management but face key limitations. Applications are disproportionately focused on disaster response, as illustrated in Figure 1, highlighting the need to expand their use across all disaster phases. Additionally, there is a gap between researchers focusing on novel architectures and practitioners needing disaster-specific solutions, necessitating LLMs tailored to various challenges.
Dataset Scarcity. Current datasets are heavily skewed toward classification tasks, leaving other critical areas underrepresented. Creating high-quality datasets requires significant effort and resources, making the process both labor- and data-intensive. Innovative solutions like synthetic data generation Kalluri et al. (2024) are promising to expand coverage across disaster-related scenarios.
Efficient Deployment. large-scale LLMs face latency and computational challenges, hindering real-time decision-making in emergent disaster scenarios. Lightweight models provide a faster, resource-efficient alternative Saleem et al. (2024), but they often sacrifice accuracy and robustness. Advancing these models to balance efficiency with reliability is crucial for effective disaster management.
Robust Generation. Decoder-based LLMs may hallucinate, where outputs can be factually inaccurate, posing serious risks in disaster contexts, such as false evacuation routes, resource misallocation, and potential loss of lives. Integrating RAG with external knowledge bases Colverd et al. (2023), extensive domain-specific training Lamsal et al. (2024a), and uncertainty estimation Xu et al. can potentially deal with hallucinated outputs.
@@ 6 Conclusion
This paper surveys LLMs in disaster management across four disaster phases, with a taxonomy integrating application scenarios, specific tasks, and the architectures of models addressing these tasks. With introducing publicly available datasets and identifying challenges, we inspire the collaborative efforts between AI researchers and decision-makers, in order to fully leveraging the power of LLMs for more resilient communities and proactive disaster management practices.
@@
This work focuses exclusively on disaster management applications only where existing LLMs have been utilized, leaving out other potential scenarios that have not yet been explored by current LLM research. While these unexplored areas hold significant promise for future advancements, they are beyond the scope of this work due to space constraints. Additionally, we include only a subset of datasets employed in existing studies, prioritizing those that are easily accessible. Many datasets are either not open-sourced, have restrictive access policies, or lack assured quality, making them less viable for reproducibility and further research.
@@ A.1 Summary Table
Table LABEL:tab:model provides a summary of the surveyed papers with their disaster phases, application scenarios, specific tasks, and architecture types.
@@ A.2 Pipeline Illustration
In this section, we present Figure 2 to illustrate the use of LLMs in disaster management. The figure outlines the major pipelines of three LLM architectures—encoder-based, decoder-based, and multimodal—applied across four task types in our survey: classification, extraction, estimation, and generation, offering critical insights into their mechanisms for disaster management.
@@ A.3 Statistics
To provide a comprehensive overview of the current state of LLMs in disaster management, we present statistics from the surveyed papers, highlighting a significant gap between the NLP and disaster management communities. This gap underscores the pressing need for better interdisciplinary collaboration to bridge the two fields and realize the potential of LLMs in addressing disaster-related challenges.
Figure 3 depicts the counts of publications leveraging existing LLMs and developing new frameworks, which indicates that most studies are heavily application-focused, relying on fine-tuning or prompting existing LLMs for disaster management tasks. In addition, while some efforts have yielded valuable insights, the majority of these studies are predominantly concentrated on the response phase of disaster management, with limited exploration across other disaster management scenarios. Figure 4 illustrates the publication distribution across academic venues, revealing that relatively few disaster management papers are published in NLP or AI-specific conferences or journals. These findings reflect limited engagement from the LLM research community in this domain, highlighting a need to increase awareness and focus within the field.
@@ Appendix B Datasets
Table LABEL:tab:dataset provides a summary of existing publicly available datasets. For classification tasks, we exclude datasets limited to a single disaster type that are already incorporated into comprehensive benchmarks such as CrisisBench Alam et al. (2021b).
@@ B.1 Classification Datasets
  • CrisisNLP Imran et al. (2016): This dataset is a large-scale dataset including classes from humanitarian disaster responses and classes related to health emergencies. It is collected from 19 different disaster events that happened between 2013 and 2015.
  • SWDM2013 Imran et al. (2013): This dataset is utilized for relevance classification consisting of tweets from two events: (i) the Joplin collection contains tweets from the tornado that struck Joplin, Missouri on May 22, 2011; (ii) The Sandy collection contains tweets collected from Hurricane Sandy that hit Northeastern US on Oct 29, 2012.
@@
  • Eyewitness Messages Zahra et al. (2020): This dataset is proposed to identity disaster eyewitness related tweets and classifies them into three types: direct eyewitnesses, indirect eyewitnesses, and vulnerable eyewitnesses where people anticipate a disaster and were present in the region for which disaster warnings were issued. It is comprised of $14k$ data from earthquakes, hurricanes, and wildfires.
@@
Figure 3: Counts of publications leveraging existing LLMs and developing new models across four phases.
@@ B.2 Estimation Datasets
  • Did You Feel It (DYFI) Atkinson and Wald (2007): This dataset contains ground shaking intensity and geographic distribution information, collected from post-earthquake reports (through $750k$ online questionnaire responses) from individuals who experienced the event.
  • FloodDepth Akinboyewa et al. (2024): This dataset contains 150 flood photos collected online, which is utilized to estimate floodwater depth based on different reference objects, including stop signs, vehicles, and humans.
  • Behavioral Facilitation (BF) Yamamoto et al. (2022): This datasets, collected after the 2018 Hokkaido earthquake, contains data labeled with usefulness ratings with respect to behavioral facilitation information.
@@ B.3 Extraction Datasets
  • Ma et al. (2023): This dataset is used for entity and relation extraction, containing $5,560$ annotated data, $12,980$ entities, and $6,895$ relations, from report during geological hazards.
  • DisasterMM Andreadis et al. (2022): This dataset is collected from Twitter by searching for flood-related keywords. It comprises two subsets: RCTP, containing 6,672 tweets for relevance classification; LETT, consisting of 4,992 tweets used for location extraction, where words are labeled with “B-LOC” for the first word of a sequence that refers to a location, “I-LOC” for the subsequent word of a sequence that refers to a location, and “O” for any non-location word.
  • Suwaileh et al. (2022): This dataset contain $22k$ crisis-related tweets from several disasters including floods, earthquakes, and hurricane. It is labeled with location-related tags such as "inLOC" and "outLOC".
@@
  • Nguyen and Rudra (2022a): This dataset contains tweet data with annotated rationales from 4 subsets of CrisisNLP. It is utilized for rationale extraction, and the extracted rationales can assist in disaster classification.
@@ B.4 Generation Datasets
  • Vitiugin and Castillo (2022): This dataset is utilized to generate summaries for various disaster events, where the ground truth is the official report of each event.
  • CrisisFACTS McCreadie and Buntain (2023): This dataset is a multiple-stream dataset with data related to 8 crisis events from various platforms. It is utilized to consume daily multi-platform streams and produce summaries for a given information need, such as "Have airports closed?".
@@
  • FFD-IQA Sun et al. (2023): This dataset includes 2,058 images and 22,422 question-meta ground truth pairs related to the safety of individuals trapped in disaster sites and the availability of emergency services. In total, three types of questions are included: free-form questions, multiple choice questions, and yes-no questions.
  • FloodNet Rahnemoonfar et al. (2021): The dataset contains 4,500 question-image pairs collected after Hurricane Harvey. Questions are related to the building, road, and the entire image, and are divided into 4 groups: "Simple Counting", "Complex Counting", "yes/no", and "Condition Recognition".
@@
Ranade et al. (2021) Response Disaster Identification Relevance Classification Encoder No
@@
Jain et al. (2019) Response Disaster Identification Relevance Classification Encoder No

### v2

@@ 5 Challenges and Opportunities
Large Language Models (LLMs) hold great promise for disaster management but face several key limitations. Most studies deploy generic LLMs as universal solutions, overlooking domain-specific challenges and the need for tailored frameworks, as shown in Figure 3. Additionally, current applications are heavily concentrated on disaster response, leaving other phases underexplored, as illustrated in Figure 1. To fully harness the potential of LLMs in disaster management, researchers must address the disaster-specific challenges outlined below.
Dataset Construction. Current datasets are heavily skewed toward classification tasks, leaving other areas underexplored Proma et al. (2022). Additionally, raw disaster data often contains uncertainty and bias Smith and Katz (2013), posing challenges in constructing reliable datasets. Innovative approaches, such as synthetic data generation Kalluri et al. (2024), offer a promising solution to enhance dataset coverage across diverse disaster scenarios.
Efficient Deployment. Large-scale LLMs face efficiency challenges Ramesh Raja et al. (2024), limiting their viability for real-time decision-making in emergency disaster scenarios. While lightweight models offer a more efficient alternative Saleem et al. (2024), they often compromise robustness in disaster-related tasks. Developing models that balance efficiency and reliability is essential for effective disaster management.
Robust Generation. (Encoder-)decoder LLMs are prone to hallucination, generating factually inaccurate outputs that pose serious risks in disaster contexts, such as false evacuation routes, resource misallocation, and potential loss of lives. To mitigate these risks, strategies such as integrating RAG with external knowledge bases Colverd et al. (2023), domain-specific training Lamsal et al. (2024a), and uncertainty estimation Xu et al. can help reduce hallucinated outputs and improve reliability.
Unified Evaluation. Most generative benchmarks for disaster (e.g. report/summary generation) rely on reference sets produced by domain experts. While these curated answers provide high-quality supervision, the underlying annotation criteria such as what counts as a correct answer often differ from one dataset to another. Consequently, published results are difficult to compare directly, because each study is implicitly tied to its own expert standard. As a result, it’s important to build up a unified evaluation protocol to make more reliable comparisons.
@@ 6 Conclusion
This paper surveys the application of LLMs in disaster management across the four disaster phases, introducing a taxonomy that integrates application scenarios, specific tasks, and the architectures of models addressing these tasks. By presenting publicly available datasets and identifying key challenges, we aim to inspire collaborative efforts between AI researchers and decision-makers, ultimately enabling the full potential of LLMs to build more resilient communities and advance proactive disaster management practices.
Acknowledgment
This work is supported in part by the National Science Foundation (NSF) under grants IIS-2006844, IIS-2144209, IIS-2223769, CNS-2154962, BCS-2228534, BCS-2228533, CMMI-2411248, CMMI-2125326, and CMMI-2402438. All authors would like to thank the reviewers and chairs for their constructive feedback and suggestions.
@@ Acknowledgment
Survey Scope. This work focuses exclusively on disaster management applications only where existing LLMs have been utilized, leaving out other potential scenarios (e.g. repair cost evaluation during the recovery phase) that have yet to be explored in current research. In addition, we focus mostly on the natural disaster instead of man-made disaster. While these unexplored areas hold significant promise for future advancements, they fall beyond the scope of this study due to space constraints.
Categorization. This work categorizes papers based on their model architecture. However, it would be also beneficial to analyze existing papers from the other perspectives such as model size, inference efficiency, and model performance.
Datasets. Additionally, we include only a subset of datasets used in existing studies, prioritizing those that are easily accessible. Many datasets either are not open-sourced, have restrictive access policies, or lack assured quality, making them less suitable for reproducibility and further research.
@@ A.1 Summary Table
Table LABEL:tab:model summarizes the surveyed papers, detailing their disaster phases, application scenarios, specific tasks, and architecture types.
@@ A.2 Pipeline Illustration
In this section, we present Figure 4, which illustrates the role of LLMs in disaster management. The figure outlines the major pipelines of three LLM architectures—encoder-based, (encoder-)decoder, and multimodal—applied across the four task types covered in this survey: classification, extraction, estimation, and generation. This visualization provides key insights into their mechanisms and applications in disaster management.
@@ A.3 Statistics
To provide a comprehensive overview of the current state of LLMs in disaster management, we present statistics from the surveyed papers, highlighting a significant gap between the NLP and disaster management communities. This gap underscores the urgent need for stronger interdisciplinary collaboration to bridge these fields and fully harness the potential of LLMs in addressing disaster-related challenges.
Figure 3 illustrates the number of publications leveraging existing LLMs versus those developing new frameworks, revealing that most studies are heavily application-focused. The majority rely on fine-tuning or prompting existing LLMs for disaster management tasks, rather than designing novel architectures. While some efforts have provided valuable insights, most research remains concentrated on the response phase, with limited exploration across other critical disaster management scenarios. Figure 5 illustrates the distribution of publications across academic venues, revealing that relatively few disaster management papers appear in NLP- or AI-specific conferences and journals. This trend reflects limited engagement from the LLM research community in this domain, underscoring the need to increase awareness and foster greater collaboration within the field.
@@ Appendix B Datasets
Table LABEL:tab:dataset summarizes existing publicly available datasets. For classification tasks, we exclude datasets that focus on a single disaster type if they are already incorporated into comprehensive benchmarks such as CrisisBench Alam et al. (2021b).
@@ B.1 Classification Datasets
  • CrisisNLP Imran et al. (2016): This dataset is a large-scale dataset that includes classes from humanitarian disaster responses and classes related to health emergencies. It is collected from 19 different disaster events that happened between 2013 and 2015.
  • SWDM2013 Imran et al. (2013): This dataset is utilized for relevance classification that consists of tweets from two events: (i) the Joplin collection contains tweets from the tornado that struck Joplin, Missouri on May 22, 2011; (ii) The Sandy collection contains tweets collected from Hurricane Sandy that struck the Northeastern US on Oct 29, 2012.
@@
  • Eyewitness Messages Zahra et al. (2020): This dataset is designed to identify disaster eyewitness-related tweets and classify them into three categories: direct eyewitnesses, indirect eyewitnesses, and vulnerable eyewitnesses—individuals who anticipate a disaster and are present in regions where disaster warnings have been issued. It comprises 14,000 tweets collected from earthquakes, hurricanes, and wildfires.
@@ B.2 Estimation Datasets
  • Did You Feel It (DYFI) Atkinson and Wald (2007): This dataset includes ground shaking intensity and geographic distribution information, collected from post-earthquake reports through 750,000 online questionnaire responses from individuals who experienced the event.
  • FloodDepth Akinboyewa et al. (2024): This dataset consists of 150 flood photos collected online, used to estimate floodwater depth based on various reference objects, including stop signs, vehicles, and humans.
  • Behavioral Facilitation (BF) Yamamoto et al. (2022): This dataset, collected after the 2018 Hokkaido earthquake, includes data labeled with usefulness ratings based on behavioral facilitation information.
@@ B.3 Extraction Datasets
  • Ma et al. (2023): This dataset is designed for entity and relation extraction, comprising 5,560 annotated instances, 12,980 entities, and 6,895 relations derived from reports on geological hazards.
  • DisasterMM Andreadis et al. (2022): This dataset was collected from Twitter by searching for flood-related keywords. It consists of two subsets: RCTP, which includes 6,672 tweets for relevance classification, and LETT, which contains 4,992 tweets used for location extraction. In the LETT subset, words are annotated with "B-LOC" for the first word in a sequence referring to a location, "I-LOC" for subsequent words within the same location sequence, and "O" for words that do not correspond to a location.
  • Suwaileh et al. (2022): This dataset contains 22,000 crisis-related tweets from various disasters, including floods, earthquakes, and hurricanes. It is annotated with location-related tags such as "inLOC" and "outLOC."
@@
  • Nguyen and Rudra (2022a): This dataset contains tweet data with annotated rationales from 4 subsets of CrisisNLP. It is used for rationale extraction, and the extracted rationales can assist in disaster classification.
@@ B.4 Generation Datasets
  • Vitiugin and Castillo (2022): This dataset is used to generate summaries of various disaster events, with the official report of each event serving as the ground truth.
  • CrisisFACTS McCreadie and Buntain (2023): This dataset is a multi-stream collection comprising data from eight crisis events gathered across various platforms. It is designed to process daily multi-platform streams and generate summaries based on specific information needs, such as "Have airports closed?"
@@
  • FFD-IQA Sun et al. (2023): This dataset comprises 2,058 images and 22,422 question-meta ground truth pairs related to the safety of individuals trapped in disaster sites and the availability of emergency services. It includes three types of questions: free-form, multiple-choice, and yes-no questions.
  • FloodNet Rahnemoonfar et al. (2021): This dataset consists of 4,500 question-image pairs collected after Hurricane Harvey. The questions pertain to buildings, roads, and entire scenes, categorized into four groups: "Simple Counting," "Complex Counting," "Yes/No," and "Condition Recognition."
@@
Dar et al. (2025) Response Disaster Information Coordination Information Classification Encoder Yes